OpenStudy (august899):

Identify the coterminal angle between 0 and 2π. Then evaluate the function:Cos 19π/3

jimthompson5910 (jim_thompson5910):

Step 1) use a calculator to compute the approximate value of `19pi/3`. Is the result between 0 and 6.28 ? If so, then the angle is already between 0 and 2pi. If not, then jump to step 2 step 2) If the angle is larger than 6.28, then subtract off 2pi. Then jump back to step 1 (but instead of 19pi/3, it would be this new angle)

jimthompson5910 (jim_thompson5910):

hopefully that makes sense?

OpenStudy (august899):

Okay.So I did 19/3 and got 6.3 then subtracted 6.28 and got 0.02,I'm confused on what to do next.

jimthompson5910 (jim_thompson5910):

19pi/3 = 19.8967534727353 is not between 0 and 6.28 so we subtract off 2pi (19pi/3) - (2pi) (19pi/3) - (6pi/3) (19pi - 6pi)/3 13pi/3 ---------------- so the angles 19pi/3 and 13pi/3 are coterminal angles

jimthompson5910 (jim_thompson5910):

is 13pi/3 between 0 and 2pi? let's find out 13pi/3 = 13.6135681655558 so the answer is no we have to subtract off 2pi again we keep this going until we find an angle between 0 and 2pi (6.28)

jimthompson5910 (jim_thompson5910):

what is `(13pi/3) - (2pi)` equal to in terms of pi?

OpenStudy (august899):

7.3π?

OpenStudy (august899):

13pi/3=13.6-2pi=7.32

jimthompson5910 (jim_thompson5910):

\[\Large \frac{13\pi}{3} - 2\pi\] \[\Large \frac{13\pi}{3}-\frac{6\pi}{3}\] \[\Large \frac{13\pi-6\pi}{3}\] \[\Large \frac{7\pi}{3}\] So 13pi/3 and 7pi/3 are coterminal. Hopefully you agree with the steps?

OpenStudy (august899):

Ohhhh yes I understand it. Thank you so much for the help!

jimthompson5910 (jim_thompson5910):

no problem